when working alone, Gail can finish her paper route in 6hrs. When Mike substitutes for her, he can complete the job in 9 hrs. If they work together, how long should it take them ?
I dont quite get what does "paper route" mean? what is that?

Answer by venugopalramana(3286)

SEE THE FOLLOWING AND TRY
Pump A can fill a tank in 6 hours. Pump A and B together can fill the same tank in 3.5 hours. How long will it take to pump B to fill the tank alone? Can you show me how to do this problem? Thank you.
1 solutions
Answer 20208 by venugopalramana(1495)on 2006-04-17 01:14:18 
Pump A can fill a tank in 6 hours.
A FILLS IN 1 HOUR..............1/6 TANK
Pump A and B together can fill the same tank in 3.5 hours.
A&B FILL IN 1 HOUR..........1/3.5 TANK=10/35=2/7 TANK
SO IN 1 HOUR B FILLS 2/7 - 1/6 = (2*6-1*7)/(6*7)=5/42...TANK
SO PUMP B ALONE WILL TAKE 1/(5/42) HRS TO FILL THE TANK =42/5=8.4 HRS.
How long will it take to pump B to fill the tank alone?
Can you show me how to do this problem? Thank you.